OpenStudy (anonymous):

solve the following equation. Check your answer and make sure list any restrictions on x. 2/a=4/a^2

OpenStudy (anonymous):

eliminating the fraction, I got (2a^2)/a =4 so a=4/2 a=2.

OpenStudy (anonymous):

thannk u!

OpenStudy (r):

\[x \neq0\]

OpenStudy (anonymous):

what are the restrictions on x? is it x cannot = 0?

OpenStudy (anonymous):

if you set x=0, what you get is something/0 = indetermination

OpenStudy (anonymous):

sooo....what are the restrictions? lol

OpenStudy (anonymous):

x not equal to zero. That is the only restriction